OpenStudy (volpina):

A spherical balloon is being inflated at a rate of 3 cm/sec. If the volume of the balloon is 0 at time 0,at what rate is the volume increase after 5 min? Plz help, i really don't know how to start...

zepdrix (zepdrix):

Volume of a sphere is given by \(\large\rm V=\frac43\pi r^3\) And we're given... the change in volume. Ya I think that's what they mean by "inflated", air is going in at this rate. \(\large\rm V'=3\)

zepdrix (zepdrix):

So ummm.. hmm what can we do >.< Thinking...

OpenStudy (danjs):

This is being inflated so that the diameter of the balloon is increasing at a steady rate. The rate of inflation is given in a change of the diameter 3 cm/s. or also the radius increases at half that, dr/dt = 1.5 cm/s

OpenStudy (danjs):

differentiate both sides of the volume of a sphere with respect to time t \[\frac{ dV }{ dt }=\frac{ 4 }{ 3 }\pi*3r^2*\frac{ dr }{ dt }\] use the given dr/dt, and you have an equation for volume at some time t